Flashcards
Module 3 Flashcards
Cards in this deck (19)
- 1random experimentis a process or a course of action that results in one of a number of possible outcomes. The outcome that occurs cannot be predicted with certainty
- 2Probabilityis the likelihood that an outcome will occur as a result of a random experiment
- 3An outcomeAn outcome is defined as the result of a single trial in a random experiment; the text refers to this as a simple event.
- 4An eventis defined as the set of one or more outcomes (simple events) of a random experiment.
- 5complementof an event A is the set of all outcomes not included in the outcomes of event A; in other words, it is everything that is “not A.”
For example, if you are interested in the probability of selecting a red card from a deck of cards, the complement to this event is the probability of not selecting a red card. The sum of the probabilities associated with an event and its complement is always one - 6Mutually exclusive eventscannot occur together. They are often thought of in association with a single experiment
- 7independentEvents are independent if the occurrence of one has no bearing on the probability of the other one occurring
- 8joint probabilityis the probability of the intersection
- 9marginal probabilitycomputed by adding across rows or down columns (totals are in the margins of the table)
- 10conditional probabilitythe probability of one event given another
- 11unionthe union of A and B is the event that occurs when either A or B both occur
- 12complementis an event that occurs when event A does not occur
- 13multiplication ruleis a rule to calculate the joint probability of two events
P(A and B)=P(B)P(A|B)
independent events:
P(A and B)=P(A)P(B) - 14addition ruleto calculate the probability of the union of two events
P(A and B) = P(A)+P(B) - P(A and B)
addition rule for mutually exclusive events:
P(A or B) = P(A)+P(B) - 15random variableis a variable whose numerical value is determined by the outcome of an experiment. For example, the dollar value in an expense report is a random variable since its numerical value changes for each report selected.
- 16A continuous random variableis a variable that is not countable and can assume an infinite number of possible values — for example, time
- 17discrete random variableis a variable that is countable and can assume a limited number of values within a given range of values
- 18probability distributionlists all possible outcomes from an experiment and the probabilities associated with each outcome. For example, if you toss a coin twice, the probabilities associated with the four possible outcomes are as follows:
P(H, H) = 0.5 × 0.5 = 0.25
P(T, H) = 0.5 × 0.5 = 0.25
P(H, T) = 0.5 × 0.5 = 0.25
P(T, T) = 0.5 × 0.5 = 0.25 - 19binomial distributionis a discrete probability distribution with only two possible outcomes. These outcomes are often referred to as “success” and “failure.”
Information
- Description
- Professional Designation QU1 (Module 3: Fundamentals of probability) Flashcards on Module 3 Flashcards, created by jbbacik on 12/02/2015.
- Created by
- jbbacik · Feb 2015
- Updated
- Feb 19, 2016
- Views
- 2
- Rating